The perimeter of a rectangle is 100. The length of the rectangle is less than two times the width. What are the dimensions ( Length and Width ) of the rectangle?
please and thank you for your time. Also, if it's possible to show all work - I'd really appreciate it !

Answer by checkley77(12844)

P=2L+2W
L=2W
100=2*2W+2W
100=4W+2W
100=6W
W=100/6
W=16.667 ANS. FOR THE WIDTH.
L=2*16.667
L=33.333 ANS.
PROOF:
100=2*33.333+2*16.667
100=66.667+33.333
100=100
L<33.333 ANSWERS THE QUESTION AS PRESENTED.